OpenStudy (anonymous):

Simplify the expression : 7w - 4w

OpenStudy (anonymous):

3w

OpenStudy (anonymous):

\[ 7w - 4w = (7-4)w = 3w. \]

OpenStudy (anonymous):

Simplifying 7w + -4w Combine like terms: 7w + -4w = 3w 3w